OpenStudy (mahoganie.carson):

Please Help I Will Give A MEDAL !!!!! 1. John Wants To Build A Fence Around His Pool The Pool Is 22 Feet Long By 23 Feet Wide The Fence Is To Be 15 Feet From The Edge Of The Pool What Are The Outside Dimensions Of The Area Surrounding the Pool and How Many Feet Of Fencing Will John Need. A) 37 ft by 38 ft 75 ft B) 37 ft by 38 ft 150 ft C) 52 ft by 53 ft 105 ft D) 52 ft by 53 ft 210 ft

OpenStudy (janu16):

One side of fence = 22 + 15 + 15 = 52 feet The other side of fence = 23 + 15 + 15 = 53 feet The fence is thus 52 feet by 53 feet. Length of fencing needed = 52 feet + 53 feet + 52 feet + 53 feet = 210 feet

OpenStudy (mahoganie.carson):

Thank you for helping me :)

OpenStudy (janu16):

no problem:)